Estimate the number of electrons in 100g of water . How much is the total negative charge on these electrons ?

Grade:12

2 Answers

Arun Kumar IIT Delhi
9 years ago
one molecule of water has (2+8)electrons


Arun Kumar
IIT Delhi
Askiitians Faculty
vishal kumar
6 years ago
N=100/18=> no. Of molecule`s=[(100/18*6.023*10²³)]One molecule of water has (2+8)electrons Ne-=[100/18*6.023*10²³]Total charge=1.6*10-¹9 *Ne-v